OpenStudy (anonymous):

HELP WILL MEDAL AND FAN Write an equation of the direct variation that includes the point (–10, –17) y = -1/17x y = 10/17x y = 1 17/10x y = -17/10x

OpenStudy (anonymous):

If you replace x with -10, what equation will equal -17?

OpenStudy (anonymous):

B? @10115658

OpenStudy (anonymous):

(10/17) x (-10)=?

OpenStudy (anonymous):

-100/17

OpenStudy (anonymous):

that does not equal -17 so B is incorrect

OpenStudy (anonymous):

Oh no I meant D, sorry

OpenStudy (anonymous):

(-17/10) x -10 =?

OpenStudy (anonymous):

=17

OpenStudy (anonymous):

17 doen't equal -17, so D is incorrect

OpenStudy (anonymous):

I dont think C is correct..

OpenStudy (anonymous):

i'm pretty sure C is supposed to be 17/10x

OpenStudy (anonymous):

(17/10) x (-10) = -17

OpenStudy (anonymous):

therefore C is the right answer

OpenStudy (anonymous):

no the 1 is there

OpenStudy (anonymous):

is the x supposed to be in the denominator for these equations?

OpenStudy (anonymous):

no on the side

OpenStudy (anonymous):

I don't think any of these equations have that coordinate

OpenStudy (anonymous):

those are the answer choices though

OpenStudy (anonymous):

Sorry dude, i'm not sure